OpenStudy (anonymous):

3x - 9y = 3 and 6x - 3y = -24 Solve and please tell me which method you used.

OpenStudy (shamil98):

3x - 9y = 3 6x -3y = -24 let's use elimination, divide the second equation by -3. your two new equations are: 3x - 9y = 3 -3x + y = 8 combine and solve for y now.

OpenStudy (anonymous):

So I divide 8 by -3?

OpenStudy (shamil98):

you add the equations together..

OpenStudy (anonymous):

Oh Oh I see

OpenStudy (shamil98):

you could do substitution as well. 3x - 9y = 3 3x = 9y + 3 x = 3y + 1 6(3y+1) -3y = -24

OpenStudy (shamil98):

ah i made an error with the first method

OpenStudy (shamil98):

3x - 9y = 3 6x -3y = -24 you would divide by -2 not -3, my bad. 3x - 9y = 3 -3x +3/2y = 12

OpenStudy (shamil98):

anyways, just use substitution, it'll be easier..

OpenStudy (shamil98):

3x - 9y = 3 6x -3y = -24 3x - 9y = 3 x = 3y + 1 x = 3y + 1 6x -3y = -24 6(3y + 1) -3y = -24 What do you get?

OpenStudy (anonymous):

18y + 6 - 3y = -24 15y + 6 = -24 15y = -30 y = -2

OpenStudy (anonymous):

Err idk if I'm correct

OpenStudy (shamil98):

yeah you're right now plug back in y and solve for x.. x = 3y + 1

OpenStudy (anonymous):

x = -5?

OpenStudy (the_fizicx99):

It's already solved?

OpenStudy (shamil98):

yes, x = -5

OpenStudy (anonymous):

Well they want me to use 2 methods. I already used one, which is the Substitution.

OpenStudy (the_fizicx99):

@shamil98 do you want to do it? or meh?

OpenStudy (shamil98):

I already gave you the equation for elimination, just solve it..

OpenStudy (shamil98):

3x - 9y = 3 -3x +3/2y = 12 have fun

OpenStudy (anonymous):

Do I gotta combine them, correct?

OpenStudy (anonymous):

So*

OpenStudy (shamil98):

yes..

OpenStudy (the_fizicx99):

Using the substitution method?

OpenStudy (anonymous):

Elimination

OpenStudy (the_fizicx99):

You asked which is the substitution.. ?

OpenStudy (anonymous):

No no, I already solved the substitution method. That's the one Shamil helped me with. They want me to use two methods. So I was asking, if for the elimination, that I have to combine the two equations.

OpenStudy (the_fizicx99):

For elimination you eliminate either the y or x and combine the two equation

OpenStudy (anonymous):

So if I were to eliminate y, I would eliminate them from both equations?

OpenStudy (shamil98):

fizi, hold on

OpenStudy (the_fizicx99):

Yes, and combine the x and the .... kk

OpenStudy (shamil98):

ellie, please read this http://www.purplemath.com/modules/systlin5.htm and then after you are enlightened work out the problem here so we can confirm it..